OpenStudy (anonymous):

Given that sin 0 = 1/3, calculate tan 0.

OpenStudy (anonymous):

I'm going to assume you mean sinΘ. Use the Pythagorean theorem to find the adjacent side value. x^2=3^2 -1^2=8 x= √8 tanΘ = 1/√8

OpenStudy (anonymous):

Yes thank you very much

OpenStudy (anonymous):

You may want to rationalize the denominator. Just multiply the top and bottom by the √8. So the fraction looks like this: tan Θ=√8/8